In Atkins, by exception, the values of a and b in the van der Waals equation are given in L.atm units! For carbon dioxide, CO2 gas the values are a = 1.453 atm.L2.mol-2 and b = 4.29x10–2 dm3. mol–1. Convert these values to SI units: Pa.m6.mol–2 for a and m3.mol–1 for b. Remember that 1 L = 1 dm3. Show your work!
Answer #1

Given

a = 1.453 atm. L2. mol-2

Now, 1 atm = 101325 Pa

1 L = 0.001 m3

1L2 = 10-6 m6

So, a = 1.453 atm. L2. mol-2 * (101325 Pa / 1 atm) * (10-6 m6 / L2) = 0.147 Pa.m6.mol-2

b = 4.29 *10-2 dm3 .mol-1 * (0.001 m3 / 1dm3) = 4.29 *10-5 m3 .mol-1